A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a long-term government anda-2. What is the expected value and standard deviation of its rate of return? (Do not round intermediate calculations. EnterPlease show excel calculations, no handwritten answer.
Answer #1

The formula for minimum risk weights in a two stock portfolio is

Ws = [06-09 *OB* PBS]/[0+os – 2* 09*0B * PBS

So, WS = (0.162-0.38*0.16 *0.10) / ( 0.162+0.382 -2*0.16*0.38 *0.1)

= 0.01952/0.15784 =0.1237 =12.37%

and WB = 1-   WS = 1-0.1237 =0.8763=87.63%

So, portfolio invested in stock is 0.1237 and portfolio invested in Bond is 0.8763

b) The Expected value of rate of return of the portfolio is the weighted average return of its constituents

Rp = WS *RS +WB *RB  = 0.1237* 0.22 + 0.8763 *0.12 = 0.13237 or 13.24%

The standard deviation of a portfolio is given by

Op = W;*W, *0; * 0;* Pij

=sqrt (0.1237^2*0.38^2+0.8763^2*0.16^2+2 *0.38*0.16*0.1237*0.8763*0.1)

= sqrt (0.023186)

=0.152269=15.23%

=0.1523

The Expected Return is 0.1324 and the standard deviation is 0.1523